# Generalization of DNNs:

### v1

@@ Abstract
We show that deep neural networks (DNNs) can efficiently learn any composition of functions with bounded $F_{1}$-norm, which allows DNNs to break the curse of dimensionality in ways that shallow networks cannot. More specifically, we derive a generalization bound that combines a covering number argument for compositionality, and the $F_{1}$-norm (or the related Barron norm) for large width adaptivity. We show that the global minimizer of the regularized loss of DNNs can fit for example the composition of two functions $f^{*}=h\circ g$ from a small number of observations, assuming $g$ is smooth/regular and reduces the dimensionality (e.g. $g$ could be the modulo map of the symmetries of $f^{*}$), so that $h$ can be learned in spite of its low regularity. The measures of regularity we consider is the Sobolev norm with different levels of differentiability, which is well adapted to the $F_{1}$ norm. We compute scaling laws empirically and observe phase transitions depending on whether $g$ or $h$ is harder to learn, as predicted by our theory.
@@ 1 Introduction
One of the fundamental features of DNNs is their ability to generalize even when the number of neurons (and of parameters) is so large that the network could fit almost any function [48]. Actually DNNs have been observed to generalize best when the number of neurons is infinite [8, 21, 20]. The now quite generally accepted explanation to this phenomenon is that DNNs have an implicit bias coming from the training dynamic where properties of the training algorithm lead to networks that generalize well. This implicit bias is quite well understood in shallow networks [11, 37], in linear networks [24, 30], or in the NTK regime [28], but it remains ill-understood in the general deep nonlinear case.
@@
In both shallow networks and linear networks, one observes a bias towards small parameter norm (either implicit [12] or explicit in the presence of weight decay [44]). Thanks to tools such as the $F_{1}$-norm [5], or the related Barron norm [46], or more generally the representation cost [14], it is possible to describe the family of functions that can be represented by shallow networks or linear networks with a finite parameter norm. This was then leveraged to prove uniform generalization bounds (based on Rademacher complexity) over these sets [5], which depend only on the parameter norm, but not on the number of neurons or parameters.
@@
Similar bounds have been proposed for DNNs [7, 6, 40, 33, 25, 41], relying on different types of norms on the parameters of the network. But it seems pretty clear that we have not yet identified the ‘right’ complexity measure for deep networks, as there remains many issues: these bounds are typically orders of magnitude too large [29, 23], and they tend to explode as the depth $L$ grows [41].
Two families of bounds are particularly relevant to our analysis: bounds based on covering numbers which rely on the fact that one can obtain a covering of the composition of two function classes from covering of the individual classes [7, 25], and path-norm bounds which extend the techniques behind the $F_{1}$-norm bound from shallow networks to the deep case [32, 6, 23].
Another issue is the lack of approximation results to accompany these generalization bounds: many different complexity measures $R(\theta)$ on the parameters $\theta$ of DNNs have been proposed along with guarantees that the generalization gap will be small as long as $R(\theta)$ is bounded, but there are often little to no result describing families of functions that can be approximated with a bounded $R(\theta)$ norm. The situation is much clearer in shallow networks, where we know that certain Sobolev spaces can be approximated with bounded $F_{1}$-norm [5].
We will focus on approximating composition of Sobolev functions, and obtaining close to optimal rates. This is quite similar to the family of tasks considered [40], though the complexity measure we consider is quite different, and does not require sparsity of the parameters.
@@ 1.1 Contribution
We consider Accordion Networks (AccNets), which are the composition of multiple shallow networks $f_{L:1}=f_{L}\circ\cdots\circ f_{1}$, we prove a uniform generalization bound $\mathcal{L}(f_{L:1})-\tilde{\mathcal{L}}_{N}(f_{L:1})\lesssim R(\theta)\frac{\log N}{\sqrt{N}}$, for a complexity measure
$R(\theta)=\prod_{\ell=1}^{L}Lip(f_{\ell})\sum_{\ell=1}^{L}\frac{\left\|f_{\ell}\right\|_{F_{1}}}{Lip(f_{\ell})}\sqrt{d_{\ell}+d_{\ell-1}}$
that depends on the $F_{1}$-norms $\left\|f_{\ell}\right\|_{F_{1}}$ and Lipschitz constanst $Lip(f_{\ell})$ of the subnetworks, and the intermediate dimensions $d_{0},\dots,d_{L}$. This use of the $F_{1}$-norms makes this bound independent of the widths $w_{1},\dots,w_{L}$ of the subnetworks, though it does depend on the depth $L$ (it typically grows linearly in $L$ which is still better than the exponential growth often observed).
@@
We then show an approximation result: any composition of Sobolev functions $f^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ can be approximated with a network with either a bounded complexity $R(\theta)$ or a slowly growing one. Thus under certain assumptions one can show that DNNs can learn general compositions of Sobolev functions. This ability can be interpreted as DNNs being able to learn symmetries, allowing them to avoid the curse of dimensionality in settings where kernel methods or even shallow networks suffer heavily from it.
@@ 2 Accordion Neural Networks and ResNets
If we add residual connections, i.e. $f_{1:L}^{res}=(f_{L}+id)\circ\cdots\circ(f_{1}+id)$ for the same shallow nets $f_{1},\dots,f_{L}$ we recover the typical ResNets.
@@
The only difference between AccNets and FCNNs is that each weight matrix $M_{\ell}$ of the FCNN is replaced by a product of two matrices $M_{\ell}=V_{\ell}W_{\ell-1}$ in the middle of the network (such a structure has already been proposed [34, 35]). Given an AccNet one can recover an equivalent FCNN by choosing $M_{\ell}=V_{\ell}W_{\ell-1}$, $M_{0}=V_{0}$ and $M_{L+1}=W_{L}$. In the other direction there could be multiple ways to split $M_{\ell}$ into the product of two matrices, but we will focus on taking $V_{\ell}=U\sqrt{S}$ and $W_{\ell-1}=\sqrt{S}V^{T}$ for the SVD decomposition $M_{\ell}=USV^{T}$, along with the choice $d_{\ell}=\mathrm{Rank}M_{\ell}$. One can thus think of AccNets as rank-constrained FCNNs.
@@ 2.1 Learning Setup
We consider a traditional learning setup, where we want to find a function $f:\Omega\subset\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ that minimizes the population loss $\mathcal{L}(f)=\mathbb{E}_{x\sim\pi}\left[\ell(x,f(x))\right]$ for an input distribution $\pi$ and a $\rho$-Lipschitz and $\rho$-bounded loss function $\ell(x,y)\in[0,B]$. Given a training set $x_{1},\dots,x_{N}$ of size $N$ we approximate the population loss by the empirical loss ~$\tilde{\mathcal{L}}_{N}(f)=\frac{1}{N}\sum_{i=1}^{N}\ell(x_{i},f(x_{i}))$ that can be minimized.
To ensure that the empirical loss remains representative of the population loss, we will prove high probability bounds on the generalization gap ~$\tilde{\mathcal{L}}_{N}(f)-\mathcal{L}(f)$ uniformly over certain functions families $f\in\mathcal{F}$.
For regression tasks, we assume the existence of a true function $f^{*}$ and try to minimize the distance $\ell(x,y)=\left\|f^{*}(x)-y\right\|^{p}$ for $p\geq 1$. If we assume that $f^{*}(x)$ and $y$ are uniformly bounded then one can easily show that $\ell(x,y)$ is bounded and Lipschitz. We are particularly interested in the cases $p\in\{1,2\}$, with $p=2$ representing the classical MSE, and $p=1$ representing a $L_{1}$ distance. The $p=2$ case is amenable to ‘fast rates’ which take advantage of the fact that the loss increases very slowly around the optimal solution $f^{*}$, We do not prove such fast rates (even though it might be possible) so we focus on the $p=1$ case.
For classification tasks on $k$ classes, we assume the existence of a ‘true class’ function $f^{*}:\Omega\to\{1,\dots,k\}$ and want to learn a function $f:\Omega\to\mathbb{R}^{k}$ such that the largest entry of $f(x)$ is the $f^{*}(k)$-th entry. One can consider the hinge cost $\ell(x,y)=\max\{0,1-(y_{f^{*}(k)}-\max_{i\neq f^{*}(x)}y_{i})\}$, which is zero whenever the margin $y_{f^{*}(k)}-\max_{i\neq f^{*}(x)}y_{i}$ is larger than $1$ and otherwise equals $1$ minus the margin. The hinge loss is Lipschitz and bounded if we assume bounded outputs $y=f(x)$. The cross-entropy loss also fits our setup.
@@ 3 Generalization Bound for DNNs
The reason we focus on accordion networks is that there exists generalization bounds for shallow networks [5, 46], that are (to our knowledge) widely considered to be tight, which is in contrast to the deep case, where many bounds exist but no clear optimal bound has been identified. Our strategy is to extend the results for shallow nets to the composition of multiple shallow nets, i.e. AccNets. Roughly speaking, we will show that the complexity of an AccNet $f_{\theta}$ is bounded by the sum of the complexities of the shallow nets $f_{1},\dots,f_{L}$ it is made of.
We will therefore first review (and slightly adapt) the existing generalization bounds for shallow networks in terms of their so-called $F_{1}$-norm [5], and then prove a generalization bound for deep AccNets.
@@ 3.1 Shallow Networks
The complexity of a shallow net $f(x)=W\sigma(Vx+b)$, with weights $W\in\mathbb{R}^{w\times d_{out}}$ and $V\in\mathbb{R}^{d_{in}\times w}$, can be bounded in terms of the quantity $C=\sum_{i=1}^{w}\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}$. First note that the rescaled function $\frac{1}{C}f$ can be written as a convex combination $\frac{1}{C}f(x)=\sum_{i=1}^{w}\frac{\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}{C}\bar{W}_{\cdot i}\sigma(\bar{V}_{i\cdot}x+\bar{b}_{i})$ for $\bar{W}_{\cdot i}=\frac{W_{\cdot i}}{\left\|W_{\cdot i}\right\|}$, $\bar{V}_{i\cdot}=\frac{V_{i\cdot}}{\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}$, and $\bar{b}_{i}=\frac{b_{i}}{\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}$, since the coefficients $\frac{\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}{C}$ are positive and sum up to 1. Thus $f$ belongs to $C$ times the convex hull
@@
We call this the $F_{1}$-ball since it can be thought of as the unit ball w.r.t. the $F_{1}$-norm $\left\|f\right\|_{F_{1}}$ which we define as the smallest positive scalar $s$ such that [^1] This construction can be used for any convex set $B$ that is symmetric around zero ($B=-B$) to define a norm whose unit ball is $B$. This correspondence between symmetric convex sets and norms is well known. $\frac{1}{s}f\in B_{F_{1}}$. For more details in the single output case, see [5].
@@
For any input distribution $\pi$ supported on the $L_{2}$ ball $B(0,b)$ with radius $b$, we have with probability $1-\delta$, over the training samples $x_{1},\dots,x_{N}$, that for all $f\in B_{F_{1}}(0,R)=R\cdot B_{F_{1}}$
$\mathcal{L}(f)-\tilde{\mathcal{L}}_{N}(f)\leq\rho bR\sqrt{d_{in}+d_{out}}\frac{\log N}{\sqrt{N}}+c_{0}\sqrt{\frac{2\log\nicefrac{{2}}{{\delta}}}{N}}$
This theorem is a slight variation of the one found in [5]: we simply generalize it to multiple outputs, and also prove it using a covering number argument instead of a direct computation of the Rademacher complexity, which will be key to obtaining a generalization bound for the deep case. But due to this change of strategy we pay a $\log N$ cost here and in our later results. We know that the $\log N$ term can be removed in Theorem 1 by switching to a Rademacher argument, but we do not know whether it can be removed in deep nets.
Notice how this bound does not depend on the width $w$, because the $F_{1}$-norm (and the $F_{1}$-ball) themselves do not depend on the width. This matches with empirical evidence that shows that increasing the width does not hurt generalization [8, 21, 20].

### v2

# How DNNs break the Curse of Dimensionality: Compositionality and Symmetry Learning

@@ Abstract
We show that deep neural networks (DNNs) can efficiently learn any composition of functions with bounded $F_{1}$-norm, which allows DNNs to break the curse of dimensionality in ways that shallow networks cannot. More specifically, we derive a generalization bound that combines a covering number argument for compositionality, and the $F_{1}$-norm (or the related Barron norm) for large width adaptivity. We show that the global minimizer of the regularized loss of DNNs can fit for example the composition of two functions $f^{*}=h\circ g$ from a small number of observations, assuming $g$ is smooth/regular and reduces the dimensionality (e.g. $g$ could be the quotient map of the symmetries of $f^{*}$), so that $h$ can be learned in spite of its low regularity. The measures of regularity we consider is the Sobolev norm with different levels of differentiability, which is well adapted to the $F_{1}$ norm. We compute scaling laws empirically and observe phase transitions depending on whether $g$ or $h$ is harder to learn, as predicted by our theory.
@@ 1 Introduction
One of the fundamental features of DNNs is their ability to generalize even when the number of neurons (and of parameters) is so large that the network could fit almost any function (Zhang et al., 2017). Actually DNNs have been observed to generalize best when the number of neurons is infinite (Belkin et al., 2019; Geiger et al., 2019; 2020). The now quite generally accepted explanation to this phenomenon is that DNNs have an implicit bias coming from the training dynamic where properties of the training algorithm lead to networks that generalize well. This implicit bias is quite well understood in shallow networks (Chizat & Bach, 2018; Rotskoff & Vanden-Eijnden, 2018), in linear networks (Gunasekar et al., 2018; Li et al., 2020), or in the NTK regime (Jacot et al., 2018), but it remains ill-understood in the general deep nonlinear case.
@@
In both shallow networks and linear networks, one observes a bias towards small parameter norm (either implicit (Chizat & Bach, 2020) or explicit in the presence of weight decay (Wang & Jacot, 2024)). Thanks to tools such as the $F_{1}$-norm (Bach, 2017), or the related Barron norm (Weinan et al., 2019), or more generally the representation cost (Dai et al., 2021), it is possible to describe the family of functions that can be represented by shallow networks or linear networks with a finite parameter norm. This was then leveraged to prove uniform generalization bounds (based on Rademacher complexity) over these sets (Bach, 2017), which depend only on the parameter norm, not on the number of neurons or parameters.
@@
Similar bounds have been proposed for DNNs (Bartlett et al., 2017; Golowich et al., 2020; Barron & Klusowski, 2019; Schmidt-Hieber, 2020; Nitanda & Suzuki, 2020; Hsu et al., 2021; Sellke, 2024), relying on different types of norms on the parameters of the network. Analogues of these bounds have been extended to other architectures, such as CNNs (Ledent et al., 2021; Graf et al., 2022; Truong, 2022; Galanti et al., 2023). But there remains many issues: these bounds are typically orders of magnitude too large (Jiang et al., 2019; Gonon et al., 2023), and they tend to explode as the depth $L$ grows (Sellke, 2024).
It is difficult to compare these bounds, since they involve many terms that have complex behavior as the depth increases. To justify our new complexity measure $R(\theta)$ and bound, we pair it with an approximation result on composition of Sobolev functions (functions with a certain decay of their Fourier or spherical harmonics coefficients), which implies that DNNs can learn such composite functions with almost optimal rates, allowing to break the curse of dimensionality in certain cases.
Previous works has shown that DNNs with either bounded width/depth (Schmidt-Hieber, 2020) or with bounded number of non-zero parameters (Galanti et al., 2023; Poggio & Fraser, 2024) can effectively learn such compositional functions. This paper shows that these strong sparsity assumptions are not strictly necessary for DNNs to learn such compositional functions.
The family of composite Sobolev functions is also a useful theoretical baseline to compare different norm-based bounds, by checking which complexity measure lead to tighter rates. This allows us to show the importance of replacing operator norms by Lipschitz constants and other design choices.
@@ 1.1 Contribution
We consider Accordion Networks (AccNets), which are the composition $f_{L:1}=f_{L}\circ\cdots\circ f_{1}$ of multiple shallow networks $f_{\ell}(x)=W_{\ell}\sigma(V_{\ell}x+b_{\ell})$, we prove a uniform generalization bound for the MSE $\sqrt{\mathcal{L}(f_{L:1})}-\sqrt{\tilde{\mathcal{L}}_{N}(f_{L:1})}\lesssim\sqrt{\frac{R(\theta)}{\sqrt{N}}}$, for a complexity measure
$R(\theta)=\left(\prod_{\ell=1}^{L}Lip(f_{\ell})\right)\left(\sum_{\ell=1}^{L}\frac{\left\|f_{\ell}\right\|_{F_{1}}}{Lip(f_{\ell})}\sqrt{d_{\ell}+d_{\ell-1}}\right)$
that depends on the $F_{1}$-norms $\left\|f_{\ell}\right\|_{F_{1}}$ (which is upper bounded by the parameter norm of the corresponding shallow network $\|W_{\ell}\|^{2}_{F}+\|V_{\ell}\|^{2}_{F}+\|b_{\ell}\|^{2}$) and Lipschitz constanst $Lip(f_{\ell})$ of the subnetworks, and the intermediate dimensions $d_{0},\dots,d_{L}$. This use of the $F_{1}$-norms makes this bound independent of the widths $w_{1},\dots,w_{L}$ of the subnetworks, though it does depend on the depth $L$ (it typically grows linearly in $L$ which is still better than the exponential growth often observed).
@@
We then show an approximation result: any composition of Sobolev functions $f^{*}=f_{L^{*}}^{*}\circ\cdots\circ f_{1}^{*}$ can be approximated with a network with either a bounded complexity $R(\theta)$ or a slowly growing one. Thus under certain assumptions one can show that DNNs can learn general compositions of Sobolev functions (i.e. functions whose first $\nu\in\mathbb{N}$ derivatives are bounded in $L_{2}$ norm). This ability can be interpreted as DNNs being able to learn symmetries, allowing them to avoid the curse of dimensionality in settings where kernel methods or even shallow networks suffer heavily from it.
@@ 2 Accordion Neural Networks and ResNets
If we add residual connections, i.e. $f_{L:1}^{res}=(f_{L}+id)\circ\cdots\circ(f_{1}+id)$ for the same shallow nets $f_{1},\dots,f_{L}$ we recover the typical ResNets.
@@
The only difference between AccNets and FCNNs is that each weight matrix $M_{\ell}$ of the FCNN is replaced by a product of two matrices $M_{\ell}=V_{\ell}W_{\ell-1}$ in the middle of the network (such a structure has already been proposed (Ongie & Willett, 2022; Parkinson et al., 2023)). Given an AccNet one can recover an equivalent FCNN by choosing $M_{\ell}=V_{\ell}W_{\ell-1}$, $M_{0}=V_{0}$ and $M_{L+1}=W_{L}$. In the other direction there could be multiple ways to split $M_{\ell}$ into the product of two matrices, but we will focus on taking $V_{\ell}=U\sqrt{S}$ and $W_{\ell-1}=\sqrt{S}V^{T}$ for the SVD decomposition $M_{\ell}=USV^{T}$, along with the choice $d_{\ell}=\mathrm{Rank}M_{\ell}$. One can thus think of AccNets as rank-constrained FCNNs.
@@ 2.1 Learning Setup
We consider a traditional MSE regression setup, where we want to find a function $f:\Omega\subset\mathbb{R}^{d_{in}}\to\mathbb{R}^{d_{out}}$ that minimizes the population loss $\mathcal{L}(f)=\mathbb{E}_{x\sim\pi}\left\|f(x)-f^{*}(x)\right\|^{2}$ for the ‘true function’ $f^{*}$, and input distribution $\pi$. Given a training set $x_{1},\dots,x_{N}$ of size $N$ we approximate the population loss by the empirical loss ~$\tilde{\mathcal{L}}_{N}(f)=\frac{1}{N}\sum_{i=1}^{N}\left\|f(x_{i})-f^{*}(x_{i})\right\|^{2}$ that can be minimized.
To ensure that the empirical loss remains representative of the population loss, we will prove high probability bounds on the generalization gap $\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}=O(\sqrt{RN^{-\frac{1}{2}}})$ uniformly over function families with complexity bounded by $R$. Thus as long as the train error is small enough ~$\tilde{\mathcal{L}}_{N}(f)=O(RN^{-\frac{1}{2}})$, so will the test error $\mathcal{L}(f)=O(RN^{-\frac{1}{2}})$. This type of bound on the gap between squared roots of the losses is more natural for the MSE loss and allows for so-called fast rates under the assumption that the train error is itself small. Note that our generalization bounds rely on covering number arguments, so they could easily be applied to Lipschitz losses using Dudley’s theorem.
For simplicity of analysis, we will assume that the true function $f^{*}$ and estimator $\hat{f}$ are uniformly bounded $\|f^{*}\|_{\infty},\|\hat{f}\|_{\infty}\leq B$. This can be guaranteed easily by adding a renormalizing nonlinearity at the end of the network $\gamma_{B}(x)=\frac{Bx}{\max\{B,\|x\|\}}$. Since this is a contraction it does not affect the covering number argument that we rely on.
@@ 3 Generalization Bound for DNNs
The reason we focus on accordion networks is that there exists generalization bounds for shallow networks (Bach, 2017; Weinan et al., 2019), that are (to our knowledge) widely considered to be tight, which is in contrast to the deep case, where many bounds exist but no clear optimal bound has been identified. Our strategy is to extend the results for shallow nets to the composition of multiple shallow nets, i.e. AccNets. Roughly speaking, we will show that the complexity of an AccNet $f_{\theta}$ is bounded by the sum of the complexities of the shallow nets $f_{1},\dots,f_{L}$ it is made of.
We will therefore first review (and slightly adapt) the existing generalization bounds for shallow networks in terms of their so-called $F_{1}$-norm (Bach, 2017), and then prove a generalization bound for deep AccNets.
@@ 3.1 Shallow Networks
The complexity of a shallow net $f(x)=W\sigma(Vx+b)$, with weights $V\in\mathbb{R}^{w\times d_{in}}$ and $W\in\mathbb{R}^{d_{out}\times w}$, can be bounded in terms of the quantity $C=\sum_{i=1}^{w}\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}$. First note that the rescaled function $\frac{1}{C}f$ can be written as a convex combination $\frac{1}{C}f(x)=\sum_{i=1}^{w}\frac{\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}{C}\bar{W}_{\cdot i}\sigma(\bar{V}_{i\cdot}x+\bar{b}_{i})$ for $\bar{W}_{\cdot i}=\frac{W_{\cdot i}}{\left\|W_{\cdot i}\right\|}$, $\bar{V}_{i\cdot}=\frac{V_{i\cdot}}{\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}$, and $\bar{b}_{i}=\frac{b_{i}}{\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}$, since the coefficients $\frac{\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}}{C}$ are positive and sum up to 1. Thus $f$ belongs to $C$ times the convex hull
@@
This set can be thought as the unit ball w.r.t. to the $F_{1}$-norm (Bach, 2017). The $F_{1}$-norm $\left\|f\right\|_{F_{1}}$ itself can then be defined as the smallest positive scalar $s$ such that $\frac{1}{s}f\in B_{F_{1}}$. Note that by the AM-GM inequality, we have
$\left\|f\right\|_{F_{1}}\leq C=\sum_{i=1}^{w}\left\|W_{\cdot i}\right\|\sqrt{\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}}\leq\frac{1}{2}\sum_{i=1}^{w}\left\|W_{\cdot i}\right\|^{2}+\left\|V_{i\cdot}\right\|^{2}+b_{i}^{2}$
@@
For any input distribution $\pi$ supported on the $L_{2}$ ball $B(0,b)$ with radius $b$, we have with probability $1-p$, over the training samples $x_{1},\dots,x_{N}$, that for all $f\in\{f:\|f\|_{F_{1}}\leq R,\|f\|_{\infty}\leq B\}$
$\sqrt{\mathcal{L}(f)}-\sqrt{\tilde{\mathcal{L}}_{N}(f)}\leq c_{0}\sqrt{BRN^{-\frac{1}{2}}}+c_{1}B\sqrt{\frac{-\log\nicefrac{{p}}{{2}}}{N}}.$
Therefore if ~$\tilde{\mathcal{L}}_{N}(f)=O(BRN^{-\frac{1}{2}})$ then $\mathcal{L}(f)=O(BRN^{-\frac{1}{2}})$.
This theorem is a slight variation of the one found in (Bach, 2017): we simply generalize it to multiple outputs, and apply to the (non-Lipschitz) MSE loss instead of a general Lipschitz loss. The proof technique however relies on covering numbers rather than Rademacher complexity, which will be key to obtaining a generalization bound for the deep case.
Notice how this bound does not depend on the width $w$, because the $F_{1}$-norm (and the $F_{1}$-ball) themselves do not depend on the width. This matches with empirical evidence that shows that increasing the width does not hurt generalization (Belkin et al., 2019; Geiger et al., 2019; 2020).